A piston/cylinder with a cross- sectional area of .01 m^2 has a piston mass of 100 kg resting on stops. With an outside atmospheric pressure of 120 kPa , what should the water pressure be to lift the piston?
Answer #1

answer) taking the sum of the forces on the piston we have

\sumF=0

Pw=water pressure, Ap=area of piston, m=mass of piston, Po=atmospheric pressure

PwAp-PoAp-mg=0

Pw=PoAp+mg/Ap

Pw=Po+mg/Ap

Pw=120kPa+100kg*9.8m/s2/0.01m2=120kPa+98kPa=218kPa

so the answer is 218 kPa.